# The Scene Language: Representing Scenes with Programs, Words, and Embeddings

### v1

@@ C.3 Minecraft Renderer
Specifically, $\Theta$ is specified in the docstring from Section D.4 and is expanded below. We introduce two types of primitives that let us construct in-game elements.
@@
We also are able to translate easily from Minecraft renderings to Mitsuba renderings, by converting Minecraft blocks to corresponding cuboids in Mitsuba. We also color the Mitsuba blocks accordingly to the average color of the Minecraft block.
@@
Similar to Mitsuba, function parameters from LM-generated programs are directly used for rendering without CLIP encoding or parameterization.
C.4 Text-to-Image (T2I) Model Renderer
@@
We employ MIGC (Zhou et al., 2024a) as the backbone model for this renderer, which originally uses a CLIP text encoder (Radford et al., 2021) and a pre-trained UNet from Stable Diffusion (Rombach et al., 2022) for layout-conditioned text-to-image generation. The parameter space $\Theta$ for this renderer is the CLIP text embedding space.
@@
Function parameters from LM-generated programs are directly encoded by the CLIP text encoder using the language templates described in Section 5.
Appendix D Language Model Prompts
D.1 Text- and Image-Conditioned Scene Generation
@@
D.2 Scene Editing
For scene editing (Section 6.2), we prompt the LM in two rounds, first with a textual scene description with the same protocol from Section 6.1, and then with an editing instruction, e.g., "move the apple to the left". In the second round, the system prompt remains the same as Section D.1. The user prompt is as follows, where {program} is the LM output from first round, and {task} is the editing instruction.
@@ D.2 Scene Editing
The full user prompt for 4D generation is displayed below.
@@
D.4 Minecraft Rendering
@@
Appendix E Experiment Details
E.1 Textual Inversion Optimization
To obtain image-conditioned embedding, we follow the procedure proposed in Gal et al. (2023). For the input image $I$ and text prompt $c_{j}$, we first use $c_{j}$ as guidance of GroundingSAM to obtain the desired mask of the corresponding entity. The cropped region is pad to square and resized to desired resolution, resulting in image target $I_{j}$. The background of $I_{j}$ is set to random grayscale color as used in Shi et al. (2024).
$v_{j1}^{*},v_{j2}^{*}=\operatorname*{arg\,min}_{v_{j1},v_{j2}}\mathbb{E}_{\eta\sim\mathcal{U}(0,1),\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\hat{\epsilon}_{\text{MVDream}}(\alpha_{\eta}I_{j}+\alpha_{\eta}\epsilon,\eta,z_{j}(v_{j1},v_{j2}))\|_{2}^{2}\Big{]}.$
For each entity, we optimize the corresponding embeddings for $100$ iterations with learning rate $1$e-$2$. Empirically we find this setting is enough to fit the image conditions. After textual inversion, the embedding $z_{j}$ is computed with optimized token embeddings, and used to guide the entity optimization as explaint in Appendix C.
E.2 GraphDreamer Implementation
[…]
After generating scene graphs with aforementioned prompt, we follow the released implementation to run Graphdreamer optimization. Each object in nodes list is represented as an individual object SDF and color field, and adopts Instant-NGP (Müller et al., 2022) as the backbone. The SDS loss optimizes not only each object, but also object pairs described in edges list. Note that Graphdreamer didn’t explicitly model the repetition of objects when dealing with scenes containing repetitive components. As a result, when the raw scene graph output contains too many objects, we’ll rerun the graph generation and add ”The maximum number of objects is three.” in text prompt to avoid reaching memory limitation during optimization.
E.3 User Study Details
Appendix F Language Model Outputs
F.1 Text-Conditioned 3D Generation
@@
F.2 Text-Conditioned 3D Editing
In this task, for each example, we prompt LM in two rounds—the first round with a scene description as task specification, and the second round with editing instructions. LM outputs in the first round are omitted.
Example 1.
[…]
Example 2.
[…]
@@ F.2 Text-Conditioned 3D Editing
F.3 Text-Conditioned 4D Generation

### v2

@@ B.2.3 Minecraft Renderer
Specifically, $\Theta$ is specified in the docstring from Sec. F.1.4 and is expanded below. We introduce two types of primitives that let us construct in-game elements.
@@
Similar to Mitsuba, function parameters from LM-generated programs are directly used for rendering without CLIP encoding or reparameterization.
B.2.4 Text-to-Image (T2I) Model Renderer
@@
We employ MIGC [42] as the backbone model for this renderer, which originally uses a CLIP text encoder [26] and a pre-trained UNet from Stable Diffusion [28] for layout-conditioned text-to-image generation. The parameter space $\Theta$ for this renderer is the CLIP text embedding space.
@@
Function parameters from LM-generated programs are directly encoded by the CLIP text encoder using the language templates described in Sec. 5.
Appendix C Experiment Details
C.1 Computation Cost
On one scene, LM inference takes $<$1 min, primitives-based rendering takes $<$1 min, SDS-based rendering takes $\sim$30 min/object. All experiments run on 1 A5000 GPU with 48GB memory.
C.2 Textual Inversion Optimization
To obtain image-conditioned embedding, we follow the procedure proposed in [6]. For the input image $I$ and text prompt $c_{j}$, we first use $c_{j}$ as guidance of GroundingSAM [27] to obtain the desired mask of the corresponding entity. The cropped region is pad to square and resized to desired resolution, resulting in image target $I_{j}$. The background of $I_{j}$ is set to random grayscale color as used in [31].
$v_{j1}^{*},v_{j2}^{*}=$
$\operatorname*{arg\,min}_{v_{j1},v_{j2}}\mathbb{E}_{\eta,\epsilon}\|\epsilon-\hat{\epsilon}_{\text{MVDream}}(\alpha_{\eta}I_{j}+\alpha_{\eta}\epsilon,\eta,z_{j}(v_{j1},v_{j2}))\|_{2}^{2}.$
For each entity, we optimize the corresponding embeddings for $100$ iterations with learning rate $1$e-$2$. Empirically we find this setting is enough to fit the image conditions. After textual inversion, the embedding $z_{j}$ is computed with optimized token embeddings, and used to guide the entity optimization as explained in Sec. B.2.
C.3 GraphDreamer Implementation
[…]
After generating scene graphs with aforementioned prompt, we follow the released implementation from GraphDreamer to optimize for the final 3D representation, where each object in the node list is represented as an individual object SDF and color field. SDS loss is used to optimize each object and object pairs described in the edge list. When the raw scene graph output contains too many objects and exceeds the memory limitation required in the optimization, we rerun the graph generation step and add “The maximum number of objects is three” in the text prompt and rerun the optimization.
Figure 10: Text-Prompted 3D Generation Comparisons extending Fig. 4.
Methods CLIP [13] ($\uparrow$) Dynamic [12] ($\uparrow$)
4D-fy [2] 0.352 $0.2\%$
4D-fy [2] (w/ prompt variant) 0.354 $0.6\%$
Ours 0.341 5.9%
Table 6: Evaluation for Text-Prompted 4D Generation extending Tab. 4.
Figure 11: Scene Editing with Language Instructions extending Fig. 7.
Figure 12: 4D Editing.
C.4 Evaluation Metrics
C.4.1 User Study
C.4.2 CLIP-Based Evaluation
We use OpenCLIP model for measuring image-text similarity (Tabs. 3 and 4). We use the model variant with the highest ImageNet zero-shot classification accuracy, OpenCLIP-ViT-H-14-378-quickgelu.
Figure 13: Ablations.
Figure 14: Visualizations of Discriminative Maps.
Figure 15: Failure Cases.
Appendix D Extended Experiment Results
D.1 Text-Prompted 3D Scene Generation.
Qualitative examples for numeric scenes and generic scenes are shown on the website and in Fig. 10, respectively.
D.2 Text-Prompted 4D Scene Generation
Representation.
The 4D entity function mentioned in Sec. 6.1 is defined as <4D-entity-func> in the DSL defined in Tab. 5.
Analysis.
Allowing for a flexible set of primitive entities is crucial to make our representation suitable for generating 4D scenes of different scales, including objects with moving parts (e.g., the wind turbine from Fig. 6) and scenes with moving objects (e.g., the carousel). This is in contrast to prior works using primitives chosen from a fixed set [36, 30] or fixing the granularity to be object-centric [15].
Moreover, the hierarchical scene structure encapsulated by our program-based representation makes it possible to represent 4D scenes compactly, serving as a regularization for generation output. Multiple entities, such as outputs from the function horse from the carousel scene in Fig. 6, can be grouped into one function horses and thereby share the same temporal transformation. Writing composible functions for entity grouping effectively reduces the dimension of the temporal motion space and improves motion fidelity.
Extended Quantitative Results.
On top of the evaluation in Tab. 4 with the same set of prompts for 4D-fy and ours, we evaluate 4D-fy with prompt variants that more explicitly indicate scene dynamics. Specifically, we use the following three prompts: "a solar system model with moving planets", "a toy wind turbine with rotating blades", and "carousel with a small canopy rotating". Results are reported as 4D-fy (w/ prompt variants) in Tab. 6. We observe a small increment in CLIP similarity and dynamic degrees in 4D-fy results with the prompt variants, but still with a relatively small motion compared to ours.
D.3 Text-Prompted 3D and 4D Scene Editing
Fig. 11 shows further qualitative experiment results under the same setup as Fig. 7. The same method applies to 4D editing, with results included in Fig. 12.
D.4 Roles of Representation Components
Below we discuss the roles of representation components, $P$ (programs), $W$ (words), and $Z$ (embeddings).
We ablate the roles of $P$ and $W$ on the 3D editing task as follows: (i) No-$P$, replacing the scene language with a leaf entity list (i.e., a list of word-pose tuples) before querying language models (LMs) to edit, (ii) No-$W$, replacing function names with random strings, and (iii) No-$P$-No-$W$, removing words from (i).
We run these variants, together with baseline methods, to edit the top-left scene in Fig. 10. Input prompts and results are shown in Fig. 13(a). Removing $P$ and/or $W$ degrades the accuracy of the results, suggesting that both components contribute to an intuitive and accurate editing interface. On the other hand, the two baseline methods do not encode precise spatial information of scene components and require regenerating the scene (we use prompt “A basket upside-down with toys in the basket”), failing to preserve the original generated content.
Embeddings $Z$ are necessary for image-prompted editing (Fig. 8). For generation tasks, $W$ and $Z$ are both required by neural renderers, and we ablate $P$ by querying LMs to directly generate entity lists. As shown in Fig. 13(b), removing programs harms LM accuracy; alternative backbone Llama has inferior performance.
D.5 Visualization of Discriminative Information
Several pieces of discriminative information can be directly obtained with the proposed Scene Language: semantic maps in Fig. 14(b), as words represent per-entity semantics; instance segmentation in Fig. 14(c), as the representation is compositional with separable instances; correspondence of the repeated instances in Fig. 14(d), as programs specify repetitions existing in a scene; dense temporal correspondence for 4D scenes, as shown in Fig. 6.
Appendix E Limitations
The current inference pipeline inherits limitations of the backbone pre-trained models.
LM/VLM errors.
For text-conditioned tasks, minor variations in textual scene descriptions can lead to large quality differences in the output as shown in Fig. 15 (a). For image-conditioned tasks, input images are parsed with the backbone visual language model. In the example in Fig. 15 (b), with the same input image, parsing results have high variance across multiple inference runs.
T2I Model Bias.
While the non-neural engines adopted are fully controllable and correct, neural-based engines may introduce additional errors, e.g., the rendered umbrella is not fully folded (Fig. 10) due to the bias in T2I models. Furthermore, the controls over texture and shapes are not disentangled for neural engines, resulting in mixed texture and shape changes in Fig. 8 as they are both affected by embeddings.
While this work provides a viable inference method for the proposed representation that leverages the commonsense knowledge and code-writing capability of LMs and expressive renderings from T2I models, addressing the weaknesses inherited from these pre-trained models would further improve the robustness, complexity, and output quality for downstream tasks. We leave these as future directions to improve the inference of the Scene Language.
Appendix F Language Model Prompts and Responses
F.1 Language Model Prompts
F.1.1 Text- and Image-Conditioned Scene Generation
@@
F.1.2 Scene Editing
For text-prompted scene editing (Sec. 6.1), we prompt the LM in two rounds, first with a textual scene description with the same protocol from Sec. 6.1, and then with an editing instruction, e.g., "move the apple to the left". In the second round, the system prompt remains the same as Sec. F.1.1. The user prompt is as follows, where {program} is the LM output from first round, and {task} is the editing instruction.
@@ F.1.2 Scene Editing
The full user prompt for 4D generation is displayed below:
@@
F.1.4 Minecraft Rendering
We adapt the prompts from Sec. F.1.1 to prompt LM to generate Minecraft-compatible outputs. Note that programs generated in this section can still be rendered using Mitsuba renderer, as shown in Fig. 5, by converting each Minecraft block from LM outputs to a Mitsuba cube primitive with the corresponding location and with the color set as the average color of the Minecraft block.
@@ F.1.4 Minecraft Rendering
F.2 Language Model Outputs
F.2.1 Text-Conditioned 3D Generation
@@
F.2.2 Text-Conditioned 3D Editing
In this task, for each example, we prompt LM in two rounds—the first round with a scene description as task specification, and the second round with editing instructions. LM outputs in the first round are omitted. Rendering outputs are shown in Fig. 7 and Fig. 11.
@@ F.2.2 Text-Conditioned 3D Editing
Example 4.
[…]
Example 5.
[…]
Example 6.
Input scene description: "7 different sized Russian nesting dolls lined up". Input editing instruction: "Shuffle order and rearrange to a small semicircle".
[…]
F.2.3 Text-Conditioned 4D Generation